I need help on this word problem setting up the equation. A 2000L tank containing 550L of water is being filled with water at the rate of 75L per minute from a full 1600L tank. How long will it be before the two tanks have the same amount of water? X equals time. Not sure how to set up this equation and solve.
Found 2 solutions by rapaljer, scott8148:
Answer by rapaljer(4671)

Let x = number of minutes
75 x = amount of water transferred

1600-75x = amount of water in the larger tank after x minutes draining
550 + 75x = amount of water in the tank being filled after x minutes of filling

1600-75x = 550 + 75x

Add +75x to each side.
1600=550 + 75x+75x

Subtract 550 from each side:
1050= 150x

Divide by 150
7=x

Final Answer = 7 minutes.

Check: After 7 minutes, 7*75 = 525 gallons of water transferred

550+525 = 1075 gallons in first tank
1600-525 = 1075 gallons in second tank

R^2

Dr. Robert J. Rapalje

Answer by scott8148(6628)

water in larger tank ___ 550 + 75X

water in smaller tank __ 1600 - 75X

set them equal to find the time ___ 550 + 75X = 1600 - 75X

150X = 1050 ___ X = 7